OpenStudy (sarahc):

Find the rectangular coordinates of the point with the polar coordinates (1, 1 divided by 2 pi). (0, -1) (1, 0) (0, 1) (-1, 0)

OpenStudy (freckles):

 hmm I wonder what point that is equivalent to in Cartesian form

OpenStudy (bradely):

x = r cos theta, y = r sin theta Here, r = 1 and theta = (1/2 ) pi x = 1 cos (pi/2), y = 1 sin (pi/2) = 0, y =1 (0,1) Source: http://www.mathskey.com/question2answer/